# FlexiTex: Enhancing Texture Generation with Visual Guidance

### v1

@@ Introduction
Text-to-texture Synthesis. Generating textures on empty models is a challenging task, which requires precise alignment with geometry and semantic coherence. Early works (Chen, Yin, and Fidler 2022; Siddiqui et al. 2022; Yu et al. 2023; Gao et al. 2022; Chan et al. 2021; mitchel2024single) aim to utilize geometric prior for texture generation. They inject positional information and train geometry-aware generative models from scratch. However, these methods cannot generalize well on various categories and show blurry textures, due to the limited 3D data with high-quality textures for training.
Optimization-based methods (Poole et al. 2022; Metzer et al. 2023; Pan et al. 2024; Youwang, Oh, and Pons-Moll 2024; Chen et al. 2023b; Metzer et al. 2023; Pan et al. 2024; zhang2024dreammat) utilize Score Distillation Sampling (SDS) on pre-trained diffusion models to update textures iteratively. Inpainting-based methods (Chen et al. 2023a; Richardson et al. 2023; Cao et al. 2023; Zeng et al. 2024; Zhang et al. 2023; tang2024intex; Chen et al. 2023a; Cao et al. 2023; ahn2024contexture) design a sequential inpainting strategy, which fills blank areas of current views based on existing contents from neighbor views. Other methods (Zhang et al. 2024b; Bensadoun et al. 2024; deng2024flashtex) finetune a multi-view diffusion model to generate a $2\times 2$ gird for multi-view consistency, while they are trained on synthetic datasets and waste rich prior gained from large-scale real image datasets, showing poor diversity.
@@
Image-to-texture Synthesis. Image-to-texture methods (Richardson et al. 2023; Zeng et al. 2024; Pan et al. 2024; yeh2024texturedreamer; Chen et al. 2023b) take the image as the user input to generate textures. TEXTure (Richardson et al. 2023) finetune DreamBooth LoRA (Ruiz et al. 2022) and uses the finetuned model as the updated base model for texture inpainting. Optimization-based methods PGC-3D (Pan et al. 2024), Fantastic3D (Chen et al. 2023b) and TextureDreamer (yeh2024texturedreamer) apply SDS conditioned on input image prompts for appearance modeling. These methods require a long time for finetuning or repainting, and cannot transfer the semantic identity well to target meshes. Paint3D (Zeng et al. 2024) applies image prompts directly on UV generation in the refinement stage, but cannot generalize well on complex UV maps, usually leaving certain artifacts on final textures. Unlike these previous methods, our method can support high-quality texture generation in a training-free manner, which is flexible because it requires no additional adjustment for image prompts.
@@
Implementation Details. Our experiments are conducted on an NVIDIA A100 GPU. For denoising epoch, we use DDIM (Song, Meng, and Ermon 2020) as the sampler. We set the number of iterations to 30 steps, the CFG scale (classifier-free guidance scale) for Direction-Aware Adaptation to 12, and the scale of Visual Guidance Enhancement to 0.6. Texture warping for latent views is used in the first 24 steps. We sample 8 views for a mesh, and the elevations and azimuths are ($-180^{\circ}$, $15^{\circ}$), ($-120^{\circ}$, $-15^{\circ}$), ($-60^{\circ}$, $15^{\circ}$), ($0^{\circ}$, $-15^{\circ}$), ($60^{\circ}$, $15^{\circ}$), ($120^{\circ}$, $-15^{\circ}$), ($-180^{\circ}$, $-45^{\circ}$), ($0^{\circ}$, $45^{\circ}$). We implement the rendering function by Pytorch3D (Ravi et al. 2020; paszke2017automatic).

### v2

@@ Introduction
Text-to-texture Synthesis. Generating textures on empty models is a challenging task, which requires precise alignment with geometry and semantic coherence. Early works (Chen, Yin, and Fidler 2022; Siddiqui et al. 2022; Yu et al. 2023; Gao et al. 2022; Chan et al. 2021; Mitchel, Esteves, and Makadia 2024) aim to utilize geometric prior for texture generation. They inject positional information and train geometry-aware generative models from scratch. However, these methods cannot generalize well on various categories and show blurry textures, due to the limited 3D data with high-quality textures for training.
Optimization-based methods (Poole et al. 2022; Metzer et al. 2023; Pan et al. 2024; Youwang, Oh, and Pons-Moll 2024; Chen et al. 2023b; Metzer et al. 2023; Pan et al. 2024; Zhang et al. 2024c) utilize Score Distillation Sampling (SDS) on pre-trained diffusion models to update textures iteratively. Inpainting-based methods (Chen et al. 2023a; Richardson et al. 2023; Cao et al. 2023; Zeng et al. 2024; Zhang et al. 2023; Tang et al. 2024; Chen et al. 2023a; Cao et al. 2023; Ahn et al. 2024) design a sequential inpainting strategy, which fills blank areas of current views based on existing contents from neighbor views. Other methods (Zhang et al. 2024b; Bensadoun et al. 2024; Deng et al. 2024) finetune a multi-view diffusion model to generate a $2\times 2$ gird for multi-view consistency, while they are trained on synthetic datasets and waste rich prior gained from large-scale real image datasets, showing poor diversity.
@@
Image-to-texture Synthesis. Image-to-texture methods (Richardson et al. 2023; Zeng et al. 2024; Pan et al. 2024; Yeh et al. 2024; Chen et al. 2023b) take the image as the user input to generate textures. TEXTure (Richardson et al. 2023) finetune DreamBooth LoRA (Ruiz et al. 2022) and uses the finetuned model as the updated base model for texture inpainting. Optimization-based methods PGC-3D (Pan et al. 2024), Fantastic3D (Chen et al. 2023b) and TextureDreamer (Yeh et al. 2024) apply SDS conditioned on input image prompts for appearance modeling. These methods require a long time for finetuning or repainting, and cannot transfer the semantic identity well to target meshes. Paint3D (Zeng et al. 2024) applies image prompts directly on UV generation in the refinement stage, but cannot generalize well on complex UV maps, usually leaving certain artifacts on final textures. Unlike these previous methods, our method can support high-quality texture generation in a training-free manner, which is flexible because it requires no additional adjustment for image prompts.
@@
Implementation Details. Our experiments are conducted on an NVIDIA A100 GPU. For denoising epoch, we use DDIM (Song, Meng, and Ermon 2020) as the sampler. We set the number of iterations to 30 steps, the CFG scale (classifier-free guidance scale) for Direction-Aware Adaptation to 12, and the scale of Visual Guidance Enhancement to 0.6. Texture warping for latent views is used in the first 24 steps. We sample 8 views for a mesh, and the elevations and azimuths are ($-180^{\circ}$, $15^{\circ}$), ($-120^{\circ}$, $-15^{\circ}$), ($-60^{\circ}$, $15^{\circ}$), ($0^{\circ}$, $-15^{\circ}$), ($60^{\circ}$, $15^{\circ}$), ($120^{\circ}$, $-15^{\circ}$), ($-180^{\circ}$, $-45^{\circ}$), ($0^{\circ}$, $45^{\circ}$). We implement the rendering function by Pytorch3D (Ravi et al. 2020; Paszke et al. 2017).